Determine the amount to be paid by the buyer for full settlement of each invoice, assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period.

Merchandise Transportation Paid by Seller Transportation Terms Returns and Allowances
(a) $6,500 $301 FOB Shipping Point, 1/10, net 30 $700
(b) $8,300 $91 FOB Destination, 2/10, net 45 $1,800
a. $
b. $
Answer #1

Answers)

a) 6500 - 700 = 5800 , 5800 x 99/100 = 5742 , 5742 + 301 = $6043

b) 8300 - 1800 = 6500 , 6500 x 98/100 = $6370